# Time-dependent Regularized 13-Moment Equations with Onsager Boundary Conditions in the Linear Regime

### v1

@@ 4.3. Removal of undesired boundary layers
where $\mathbf{A}_{\mathrm{eo}}=\mathbf{A}_{\mathrm{oe}}^{\intercal}$, and $\mathbf{L}_{\mathrm{oo}}$ and $\mathbf{L}_{\mathrm{ee}}$ are symmetric and negative semidefinite matrices. To mimic the property that $\mathcal{A}_{n}$ has a non-trivial null space, we assume that the matrix $\mathbf{A}_{\mathrm{oe}}$ has rank $r$, and its singular value decomposition is
@@
where
@@
After imposing these conditions, the matrix $\tilde{\mathbf{Q}}$ is uniquely determined. Currently, we do not have theoretical guarantee of the semidefiniteness of the revised $\tilde{\mathbf{Q}}$, but this turns out to be true in all cases done in our tests.
@@ 5. One-dimensional simulations
In this section, we will apply our model to the example for one-dimensional channel flows in [16, 8]. We will show the analytic solutions of the steady-state equations of our model in section 5.1, and show the numerical solutions of our time-dependent model in section 5.2 using the finite element method.
@@ 5.1. Steady-state examples
To compare with the example in section 4.2, we consider a steady-state problem here. In fact, equations (71)–(75) can be regarded as a portion of the considering one-dimensional steady-state channel flows by the following replacement: $x\to x_{2}$, $v\to v_{2}$, $\bar{q}\to\bar{q}_{2}$ and $\bar{\sigma}\to\bar{\sigma}_{22}$. The boundary conditions (76)–(79) also share the same form of the considering boundary conditions by the above replacement together with coefficient $\hat{m}_{ij}\to{m}_{ij}$.
In the first example, we adopt the coefficients $\mathrm{Kn}=0.2$, $\theta^{W}_{l}=0$ and $\theta^{W}_{r}=0.2$, consistent with the example presented in section 4.2. The results using our final boundary conditions are depicted in figure 2, where we maintain the same notations as those used in the plots shown in figure 1 for ease of comparison. The comparison clearly demonstrates that the non-physical boundary layers have been removed with the modified boundary conditions. Furthermore, using the modified boundaries, the results from the three inverse-power-law models exhibit consistency both in regions far from the boundaries and in areas close to the boundaries.
To observe the effect of the Knudsen number, we adopt the same coefficients as the previous one except changing $\mathrm{Kn}$ to $0.4$, and plot the results for both original and fixed boundary conditions in figure 3. The effectiveness of our modified Onsager boundary conditions is again validated by the successful removal of the undesired boundary layers. Furthermore, compared to the results in figure 1, the width of the undesired boundary layers increases as the Knudsen number increases, which is consistent with our analysis of the solution in (85). Also, it can be observed that the temperature jump gets more significant as the Knudsen number increases.
Figure 3. $\bar{q}$ and $\theta$ of the one-dimensional problem (71)–(79) with two different boundary conditions. Top subfigures use coefficients $\hat{m}_{ij}$ and bottom subfigures use modified ones $m_{ij}$.
@@ 6. Conclusion
This work provides a new perspective of the derivation of time-dependent regularized 13-moment equations from linear kinetic equation under general elastic collision models. Compared with previous works such as [16] and [8], our equations possess not only the super-Burnett order, but also a symmetric structure that is useful for deriving Onsager boundary conditions. Another contribution of this work is the technique to remove undesired boundary layers, which needs to be applied since these layers are obviously unphysical. Future work includes applying this approach to nonlinear kinetic equations.
Appendix A Tables of coefficients in the inverse-power-law model
@@
$\begin{array}[]{| c | c | c | c | c | c |}\hline\cr\hline\cr\eta&m_{11}&m_{12}&m_{13}&m_{14}&m_{15}\\ 5&\frac{2}{\sqrt{2\pi}}&\frac{1}{2\sqrt{2\pi}}&\frac{8}{5\sqrt{2\pi}}&\frac{48}{25\sqrt{2\pi}}&0\\ 7&8.2699\times 10^{-1}&1.7756\times 10^{-1}&5.9524\times 10^{-1}&7.7155\times 10^{-1}&4.0454\times 10^{-2}\\ 10&8.4706\times 10^{-1}&1.6277\times 10^{-1}&5.7153\times 10^{-1}&7.7914\times 10^{-1}&6.9432\times 10^{-2}\\ 17&8.6507\times 10^{-1}&1.4961\times 10^{-1}&5.5359\times 10^{-1}&7.8843\times 10^{-1}&9.6598\times 10^{-2}\\ \infty&8.8890\times 10^{-1}&1.3234\times 10^{-1}&5.3390\times 10^{-1}&8.0443\times 10^{-1}&1.3481\times 10^{-1}\\ \hline\cr\eta&m_{21}&m_{22}&m_{23}&m_{24}&m_{25}\\ 5&-&-&-&-&-\\ 7&4.9821\times 10^{-3}&1.9210\times 10^{-2}&3.5585\times 10^{-3}&4.6126\times 10^{-3}&2.4185\times 10^{-4}\\ 10&7.5057\times 10^{-3}&3.1012\times 10^{-2}&4.9556\times 10^{-3}&6.7557\times 10^{-3}&6.0203\times 10^{-4}\\ 17&9.2216\times 10^{-3}&4.0714\times 10^{-2}&5.6679\times 10^{-3}&8.0723\times 10^{-3}&9.8900\times 10^{-4}\\ \infty&1.0730\times 10^{-2}&5.2344\times 10^{-2}&5.9826\times 10^{-3}&9.0140\times 10^{-3}&1.5106\times 10^{-3}\\ \hline\cr\eta&m_{26}&m_{27}&m_{28}\\ 5&-&-&-\\ 7&2.9445\times 10^{-2}&1.1082\times 10^{-1}&4.5197\times 10^{-4}\\ 10&5.0347\times 10^{-2}&1.9074\times 10^{-1}&9.9954\times 10^{-4}\\ 17&7.0089\times 10^{-2}&2.6780\times 10^{-1}&1.6422\times 10^{-3}\\ \infty&9.8823\times 10^{-2}&3.8322\times 10^{-1}&2.6338\times 10^{-3}\\ \hline\cr\eta&m_{31}&m_{32}&m_{33}&m_{34}&m_{35}\\ 5&\frac{1}{\sqrt{2\pi}}&\frac{1}{5\sqrt{2\pi}}&0&\frac{2}{\sqrt{2\pi}}&0\\ 7&4.0358\times 10^{-1}&7.2977\times 10^{-2}&5.8325\times 10^{-8}&7.8808\times 10^{-1}&7.6807\times 10^{-6}\\ 10&4.0655\times 10^{-1}&6.8340\times 10^{-2}&2.4645\times 10^{-7}&7.8723\times 10^{-1}&2.8477\times 10^{-5}\\ 17&4.0908\times 10^{-1}&6.4237\times 10^{-2}&5.9816\times 10^{-7}&7.9036\times 10^{-1}&6.3150\times 10^{-5}\\ \infty&4.1227\times 10^{-1}&5.8927\times 10^{-2}&1.3613\times 10^{-6}&8.0020\times 10^{-1}&1.3351\times 10^{-4}\\ \hline\cr\eta&m_{41}&m_{42}&m_{43}&m_{44}&m_{45}\\ 5&\frac{1}{\sqrt{2\pi}}&\frac{11}{5\sqrt{2\pi}}&0&\frac{2}{\sqrt{2\pi}}&0\\ 7&3.4342\times 10^{-1}&8.5528\times 10^{-1}&5.6647\times 10^{-8}&7.6541\times 10^{-1}&7.4598\times 10^{-6}\\ 10&3.0441\times 10^{-1}&8.4194\times 10^{-1}&2.3349\times 10^{-7}&7.4584\times 10^{-1}&2.6980\times 10^{-5}\\ 17&2.6826\times 10^{-1}&8.3207\times 10^{-1}&5.5187\times 10^{-7}&7.2919\times 10^{-1}&5.8263\times 10^{-5}\\ \infty&2.1746\times 10^{-1}&8.2380\times 10^{-1}&1.2025\times 10^{-6}&7.0683\times 10^{-1}&1.1793\times 10^{-4}\\ \hline\cr\eta&m_{46}&m_{47}&m_{48}\\ 5&0&0&\frac{24}{5}\\ 7&5.7974\times 10^{-1}&2.4505\times 10^{-1}&4.7008\\ 10&9.6361\times 10^{-1}&4.0914\times 10^{-1}&4.6366\\ 17&1.2954&5.5322\times 10^{-1}&4.5779\\ \infty&1.7103&7.3790\times 10^{-1}&4.4916\\ \hline\cr\hline\cr\end{array}$

### v2

@@ 4.3. Removal of undesired boundary layers
where $\mathbf{A}_{\mathrm{eo}}=\mathbf{A}_{\mathrm{oe}}^{\intercal}$, and $\mathbf{L}_{\mathrm{oo}}$ and $\mathbf{L}_{\mathrm{ee}}$ are symmetric and negative semidefinite matrices. As noted in section 4.1, the operator $\mathcal{A}_{n}$ has a non-trivial null space being a subspace of $\mathcal{P}_{\mathrm{odd}}\mathbb{V}$. Accordingly, in order to mimic this property, we assume that the matrix $\mathbf{A}_{\mathrm{oe}}$ is rank-deficient, and its singular value decomposition is
@@
where $r$ denotes the rank of $\mathbf{A}_{\mathrm{oe}}$ and
@@
After imposing these conditions, the matrix $\tilde{\mathbf{Q}}$ is uniquely determined. Currently, we do not have theoretical guarantee of the semidefiniteness of the revised $\tilde{\mathbf{Q}}$, but this turns out to be true in all cases done in our tests. The symmetry and semidefiniteness of $\tilde{\mathbf{Q}}$ indicates that the structure (6) is well maintained after the modification, and thus does not ruin the $L^{2}$ stability or the second law of thermodynamics.
To demonstrate the effect of this modification, we reconsider the example described in section 4.2, with the boundary conditions changed to the new ones introduced in this section by replacing $\hat{m}_{ij}$ with ${m}_{ij}$ in (82)–(85). The results are plotted in figure 2. A comparison with figure 1 clearly demonstrates that the non-physical boundary layers have been removed.
@@ 5. One-dimensional simulations
In this section, we will apply our model to the example for one-dimensional channel flows in [19, 10]. We will show the analytic solutions of the steady-state equations of our model and make a comparison with the results computed by direct simulation Monte Carlo (DSMC) from Bird’s code ([5]) in section 5.1. After that, the numerical solutions of our time-dependent model will be demonstrated in section 5.2 using the finite element method.
@@ 5.1. Steady-state examples
To compare with the example in sections 4.2 & 4.3, we consider a steady-state problem here. In fact, equations (77)–(81) can be regarded as a portion of the considering one-dimensional steady-state channel flows by the following replacement: $x\to x_{2}$, $v\to v_{2}$, $\bar{q}\to\bar{q}_{2}$ and $\bar{\sigma}\to\bar{\sigma}_{22}$. The boundary conditions (82)–(85) also share the same form of the considering boundary conditions by the above replacement. We adopt the coefficients $\theta^{W}_{l}=0$ and $\theta^{W}_{r}=0.2$, consistent with the example presented in section 4.2.
To compare our solutions with the DSMC results, we show the results by using the Knudsen number $\overline{\mathrm{Kn}}$ defined by the ratio of the mean free path to the characteristic length as in [5]. The relationship between $\overline{\mathrm{Kn}}$ and the parameter $\mathrm{Kn}$ in this paper is
$\mathrm{Kn}=\sqrt{\frac{\pi}{2}}\frac{15\overline{\mathrm{Kn}}}{(5-2\omega)(7-2\omega)},\qquad\omega=\frac{1}{2}+\frac{2}{\eta-1}.$
Moreover, we plot the actual heat flux $q_{2}$ instead of $\bar{q}_{2}$ (see (12)) in this example, since $q_{2}$ is obtained from the DSMC results. To observe the effect of the Knudsen number, we choose two different $\overline{\mathrm{Kn}}$ and plot the results for both original and fixed boundary conditions in figures 3 & 4. Our results are depicted as the blue, yellow, and green solid lines corresponding to $\eta=5,10$ and $\infty$. The relevant solutions obtained by DSMC are presented by dotted lines of the same colors.
Figure 3. Results of steady-state example when $\overline{\mathrm{Kn}}=0.1$. Top subfigures use coefficients $\hat{m}_{ij}$ and bottom subfigures use modified ones $m_{ij}$. DSMC solutions are given by dotted lines of the same colors.
Figure 4. Results of steady-state example when $\overline{\mathrm{Kn}}=0.05$. Top subfigures use coefficients $\hat{m}_{ij}$ and bottom subfigures use modified ones $m_{ij}$. DSMC solutions are given by dotted lines of the same colors.
The solids lines of the R13 equations generally agree with the dotted lines of the DSMC results, and this agreement is more pronounced for smaller $\mathrm{Kn}$, which validates our equations. Moreover, as the DSMC results display no boundary layer, the effectiveness of our modified Onsager boundary conditions is again validated by the successful removal of the undesired boundary layers. Furthermore, compared to the results in figure 1, the width of the undesired boundary layers decreases as the Knudsen number decreases, which is consistent with our analysis of the solution in (91). Also, it can be observed that the temperature jump gets more significant as the Knudsen number increases.
@@ 6. Conclusion
This work provides a new perspective of the derivation of time-dependent regularized 13-moment equations from linear kinetic equation under general elastic collision models. Compared with previous works such as [19] and [10], our equations possess not only the super-Burnett order, but also a symmetric structure that is useful for deriving Onsager boundary conditions. Another contribution of this work is the technique to remove undesired boundary layers, which needs to be applied since these layers are obviously unphysical.
Due to the symmetry of the equations and the Onsager boundary conditions, we expect that our equations also admit a symmetric weak form, which extends the weak form of R13 equations for Maxwell molecules ([38]). We are currently working on the finite element methods for our equations in the multidimensional case. Note that it is claimed in [27] that R13 equations can provide satisfactory predictions up to Knudsen number $0.5$, and this needs to be tested for our equations with modified boundary conditions in multidimensional applications. Meanwhile, the work on applying this approach to nonlinear kinetic equations is also ongoing.
Appendix A Galilean invariance of moment equations
To demonstrate the Galilean invariance of our moment equations, we need to prove the following invariant properties:
  • Translational invariance: given any $\boldsymbol{v}_{\text{ref}}\in\mathbb{R}^{3}$, under translation
    $\boldsymbol{x}^{\prime}=\boldsymbol{x}+t\boldsymbol{v}_{\text{ref}},\ \rho^{\prime}(\boldsymbol{x}^{\prime},t)=\rho(\boldsymbol{x},t),\ \theta^{\prime}(\boldsymbol{x}^{\prime},t)=\theta(\boldsymbol{x},t),$
    $\boldsymbol{v}^{\prime}(\boldsymbol{x},t)=\boldsymbol{v}(\boldsymbol{x}^{\prime},t)+\boldsymbol{v}_{\text{ref}},\ \boldsymbol{\bar{q}}^{\prime}(\boldsymbol{x}^{\prime},t)=\boldsymbol{\bar{q}}(\boldsymbol{x},t),\ \boldsymbol{\bar{\sigma}}^{\prime}(\boldsymbol{x}^{\prime},t)=\boldsymbol{\bar{\sigma}}(\boldsymbol{x},t),$
    equations (7)–(11) retain their form.
  • Rotational invariance: given any orthogonal matrix $\boldsymbol{R}\in\mathbb{R}^{3\times 3}$, under rotation
    $\boldsymbol{x}^{\prime}=\boldsymbol{R}\boldsymbol{x},\ \rho^{\prime}(\boldsymbol{x}^{\prime},t)=\rho(\boldsymbol{x},t),\ \theta^{\prime}(\boldsymbol{x}^{\prime},t)=\theta(\boldsymbol{x},t),$
    $\boldsymbol{v}^{\prime}(\boldsymbol{x}^{\prime},t)=\boldsymbol{R}\boldsymbol{v}(\boldsymbol{x},t),\ \boldsymbol{\bar{q}}^{\prime}(\boldsymbol{x}^{\prime},t)=\boldsymbol{R}\boldsymbol{\bar{q}}(\boldsymbol{x},t),\ \boldsymbol{\bar{\sigma}}^{\prime}(\boldsymbol{x}^{\prime},t)=\boldsymbol{R}\boldsymbol{\bar{\sigma}}(\boldsymbol{x},t)\boldsymbol{R}^{\intercal},$
    equations (7)–(11) retain their form.
Since the equations rely solely on the derivatives of $\boldsymbol{v}$ rather than on $\boldsymbol{v}$ itself, the translational invariance can be readily verified in our linear equations. Consequently, we focus on the proof of the rotational invariance in this appendix.
For (7), we have
(102) $\frac{\partial v^{\prime}_{i}}{\partial x^{\prime}_{i}}=\frac{\partial v^{\prime}_{i}}{\partial x_{k}}\frac{\partial x_{k}}{\partial x^{\prime}_{i}}=R_{ij}\frac{\partial v_{j}}{\partial x_{k}}R_{ik}=\frac{\partial v_{j}}{\partial x_{k}}\delta_{jk}=\frac{\partial v_{j}}{\partial x_{j}}$
which implies that $\nabla_{\boldsymbol{x}^{\prime}}\cdot\boldsymbol{v}^{\prime}=\nabla_{\boldsymbol{x}}\cdot\boldsymbol{v}$. Therefore, (7) retains its form under the rotational transformation.
For (8), we first calculate the derivative of $\bar{\sigma}$ as
(103) $\frac{\partial\bar{\sigma}^{\prime}_{ij}}{\partial x^{\prime}_{j}}=\frac{\partial\bar{\sigma}^{\prime}_{ij}}{\partial x_{k}}\frac{\partial x_{k}}{\partial x^{\prime}_{j}}=R_{il}R_{jl^{\prime}}\frac{\partial\bar{\sigma}_{ll^{\prime}}}{\partial x_{k}}R_{jk}=R_{il}\delta_{kl^{\prime}}\frac{\partial\bar{\sigma}_{ll^{\prime}}}{\partial x_{k}}=R_{il}\frac{\partial\bar{\sigma}_{lk}}{\partial x_{k}},$
and then we arrive at the relationship
$\frac{\partial^{2}\bar{\sigma}^{\prime}_{ij}}{\partial x^{\prime}_{i}\partial x^{\prime}_{j}}=R_{il}\frac{\partial^{2}\bar{\sigma}_{lk}}{\partial x_{k}\partial x_{k^{\prime}}}R_{ik^{\prime}}=\frac{\partial^{2}\bar{\sigma}_{lk}}{\partial x_{k}\partial x_{k^{\prime}}}\delta_{lk^{\prime}}=\frac{\partial^{2}\bar{\sigma}_{k^{\prime}k}}{\partial x_{k}\partial x_{k^{\prime}}}$
which implies that $\nabla_{\boldsymbol{x}^{\prime}}\cdot(\nabla_{\boldsymbol{x}^{\prime}}\cdot\boldsymbol{\bar{\sigma}}^{\prime})=\nabla_{\boldsymbol{x}}\cdot(\nabla_{\boldsymbol{x}}\cdot\boldsymbol{\bar{\sigma}})$. Similar to (102), we can prove that $\nabla_{\boldsymbol{x}^{\prime}}\cdot\boldsymbol{\bar{q}}^{\prime}=\nabla_{\boldsymbol{x}}\cdot\boldsymbol{\bar{q}}$. In addition, $\Delta_{\boldsymbol{x}^{\prime}}\theta^{\prime}=\Delta_{\boldsymbol{x}}\theta$ according to the rotational invariance of Laplacian. As a result, (8) is also rotational invariant.
For (9), one can easily see that
(104) $\frac{\partial\boldsymbol{v}^{\prime}}{\partial\boldsymbol{x}^{\prime}}=\boldsymbol{R}\frac{\partial\boldsymbol{v}}{\partial\boldsymbol{x}},\quad\nabla_{\boldsymbol{x}^{\prime}}\rho^{\prime}=\boldsymbol{R}\nabla_{\boldsymbol{x}}\rho,\quad\nabla_{\boldsymbol{x}^{\prime}}\theta^{\prime}=\boldsymbol{R}\nabla_{\boldsymbol{x}}\theta.$
In addition, we have $\nabla_{\boldsymbol{x}^{\prime}}\cdot\boldsymbol{\bar{\sigma}}^{\prime}=\boldsymbol{R}\nabla_{\boldsymbol{x}}\cdot\boldsymbol{\bar{\sigma}}$ by (103). For the term $\nabla\cdot(\nabla\boldsymbol{v})_{\mathrm{stf}}$, we use the fact
(105) $\begin{split}(\nabla\boldsymbol{v})_{\mathrm{stf}}=&\ \frac{1}{2}(\nabla\boldsymbol{v}+(\nabla\boldsymbol{v})^{\intercal})-\frac{1}{3}\text{tr}(\nabla\boldsymbol{v})I,\\ \nabla_{\boldsymbol{x}^{\prime}}\boldsymbol{v}^{\prime}=&\ \boldsymbol{R}\nabla_{\boldsymbol{x}}\boldsymbol{v}\boldsymbol{R}^{T},\end{split}$
and cyclic property of trace operator to show that $(\nabla_{x^{\prime}}\boldsymbol{v}^{\prime})_{\mathrm{stf}}=\boldsymbol{R}(\nabla_{x}\boldsymbol{v})_{\mathrm{stf}}\boldsymbol{R}^{T}$. Therefore, we have
$\frac{\partial}{\partial x^{\prime}_{j}}[(\nabla_{x^{\prime}}\boldsymbol{v}^{\prime})_{\mathrm{stf}}]_{ij}=R_{jk}\frac{\partial}{\partial x_{k}}R_{il}[(\nabla_{x}\boldsymbol{v})_{\mathrm{stf}}]_{ll^{\prime}}R_{jl^{\prime}}=R_{il}\frac{\partial}{\partial x_{k}}[(\nabla_{x}\boldsymbol{v})_{\mathrm{stf}}]_{lk}$
which implies that $\nabla_{\boldsymbol{x}^{\prime}}\cdot(\nabla_{\boldsymbol{x}^{\prime}}\boldsymbol{v}^{\prime})_{\mathrm{stf}}=\boldsymbol{R}\nabla_{\boldsymbol{x}}\cdot(\nabla_{\boldsymbol{x}}\boldsymbol{v})_{\mathrm{stf}}$. Similarly, we can prove such relationship for $\boldsymbol{\bar{q}}$. Combining these two equalities of $\boldsymbol{v}$ and $\boldsymbol{\bar{q}}$ in conjunction with (103) and (104), we have verified the rotational invariance of (9). The rotational invariance of (10) can be similarly proved as (9).
For (11), similar as (105), one can prove that $\nabla_{\boldsymbol{x}^{\prime}}\boldsymbol{\bar{q}}^{\prime}=\boldsymbol{R}\nabla_{\boldsymbol{x}}\boldsymbol{\bar{q}}\boldsymbol{R}^{T}$ and $(\nabla_{x^{\prime}}\boldsymbol{\bar{\sigma}}^{\prime})_{\mathrm{stf}}=\boldsymbol{R}(\nabla_{x}\boldsymbol{\bar{\sigma}})_{\mathrm{stf}}\boldsymbol{R}^{T}$. For the term $\left(\nabla(\nabla\cdot\boldsymbol{\bar{\sigma}})\right)_{\mathrm{stf}}$, it holds that
$[\nabla_{\boldsymbol{x}^{\prime}}(\nabla_{\boldsymbol{x}^{\prime}}\cdot\boldsymbol{\bar{\sigma}}^{\prime})]_{ij}=$ $\ \frac{\partial^{2}\bar{\sigma}^{\prime}_{ik}}{\partial x^{\prime}_{j}\partial x^{\prime}_{k}}$
$=$ $\ R_{jk^{\prime}}R_{il}R_{kl}R_{kl^{\prime\prime}}\frac{\partial^{2}\bar{\sigma}_{ll^{\prime}}}{\partial x_{k^{\prime}}\partial x_{l^{\prime\prime}}}=R_{jk^{\prime}}R_{il}\delta_{ll^{\prime\prime}}\frac{\partial^{2}\bar{\sigma}_{ll^{\prime}}}{\partial x_{k^{\prime}}\partial x_{l^{\prime\prime}}}=R_{il}\frac{\partial^{2}\bar{\sigma}_{ll^{\prime}}}{\partial x_{k^{\prime}}\partial x_{l}}R_{jk^{\prime}},$
which implies that $(\nabla_{\boldsymbol{x}^{\prime}}(\nabla_{\boldsymbol{x}^{\prime}}\cdot\boldsymbol{\bar{\sigma}}^{\prime}))_{\mathrm{stf}}=\boldsymbol{R}(\nabla_{\boldsymbol{x}}(\nabla_{\boldsymbol{x}}\cdot\boldsymbol{\bar{\sigma}}))_{\mathrm{stf}}\boldsymbol{R}^{\intercal}$ upon using the formula of $(\cdot)_{\mathrm{stf}}$ in (105). Furthermore, it is evident that $(\nabla^{2}_{\boldsymbol{x}^{\prime}}\theta^{\prime})_{\mathrm{stf}}=\boldsymbol{R}(\nabla^{2}_{\boldsymbol{x}}\theta)_{\mathrm{stf}}\boldsymbol{R}^{\intercal}$. Therefore, the rotational invariance of (11) has been verified.
Overall, our moment equations (7)–(11) satisfy the rotational invariance and, by encompassing the translational invariance, consequently satisfy the Galilean invariance.
Appendix B Tables of coefficients in the inverse-power-law model
@@
$\begin{array}[]{| c | c | c | c | c | c |}\hline\cr\hline\cr\eta&m_{11}&m_{12}&m_{13}&m_{14}&m_{15}\\ 5&\frac{2}{\sqrt{2\pi}}&\frac{1}{2\sqrt{2\pi}}&\frac{8}{5\sqrt{2\pi}}&\frac{48}{25\sqrt{2\pi}}&0\\ 7&8.2699\times 10^{-1}&1.7756\times 10^{-1}&5.9524\times 10^{-1}&7.7155\times 10^{-1}&4.0454\times 10^{-2}\\ 10&8.4706\times 10^{-1}&1.6277\times 10^{-1}&5.7153\times 10^{-1}&7.7914\times 10^{-1}&6.9432\times 10^{-2}\\ 17&8.6507\times 10^{-1}&1.4961\times 10^{-1}&5.5359\times 10^{-1}&7.8843\times 10^{-1}&9.6598\times 10^{-2}\\ \infty&8.8890\times 10^{-1}&1.3234\times 10^{-1}&5.3390\times 10^{-1}&8.0443\times 10^{-1}&1.3481\times 10^{-1}\\ \hline\cr\eta&m_{21}&m_{22}&m_{23}&m_{24}&m_{25}\\ 5&-&-&-&-&-\\ 7&4.9821\times 10^{-3}&1.9210\times 10^{-2}&3.5585\times 10^{-3}&4.6126\times 10^{-3}&2.4185\times 10^{-4}\\ 10&7.5057\times 10^{-3}&3.1012\times 10^{-2}&4.9556\times 10^{-3}&6.7557\times 10^{-3}&6.0203\times 10^{-4}\\ 17&9.2216\times 10^{-3}&4.0714\times 10^{-2}&5.6679\times 10^{-3}&8.0723\times 10^{-3}&9.8900\times 10^{-4}\\ \infty&1.0730\times 10^{-2}&5.2344\times 10^{-2}&5.9826\times 10^{-3}&9.0140\times 10^{-3}&1.5106\times 10^{-3}\\ \hline\cr\eta&m_{26}&m_{27}&m_{28}\\ 5&-&-&-\\ 7&2.9445\times 10^{-2}&1.1082\times 10^{-1}&4.5197\times 10^{-4}\\ 10&5.0347\times 10^{-2}&1.9074\times 10^{-1}&9.9954\times 10^{-4}\\ 17&7.0089\times 10^{-2}&2.6780\times 10^{-1}&1.6422\times 10^{-3}\\ \infty&9.8823\times 10^{-2}&3.8322\times 10^{-1}&2.6338\times 10^{-3}\\ \hline\cr\eta&m_{31}&m_{32}&m_{33}&m_{34}&m_{35}\\ 5&\frac{1}{\sqrt{2\pi}}&\frac{1}{5\sqrt{2\pi}}&0&\frac{2}{\sqrt{2\pi}}&0\\ 7&4.0358\times 10^{-1}&7.2977\times 10^{-2}&5.8325\times 10^{-8}&7.8808\times 10^{-1}&7.6807\times 10^{-6}\\ 10&4.0655\times 10^{-1}&6.8340\times 10^{-2}&2.4645\times 10^{-7}&7.8723\times 10^{-1}&2.8477\times 10^{-5}\\ 17&4.0908\times 10^{-1}&6.4237\times 10^{-2}&5.9816\times 10^{-7}&7.9036\times 10^{-1}&6.3150\times 10^{-5}\\ \infty&4.1227\times 10^{-1}&5.8927\times 10^{-2}&1.3613\times 10^{-6}&8.0020\times 10^{-1}&1.3351\times 10^{-4}\\ \hline\cr\eta&m_{41}&m_{42}&m_{43}&m_{44}&m_{45}\\ 5&\frac{1}{\sqrt{2\pi}}&\frac{11}{5\sqrt{2\pi}}&0&\frac{2}{\sqrt{2\pi}}&0\\ 7&3.4342\times 10^{-1}&8.5528\times 10^{-1}&5.6647\times 10^{-8}&7.6541\times 10^{-1}&7.4598\times 10^{-6}\\ 10&3.0441\times 10^{-1}&8.4194\times 10^{-1}&2.3349\times 10^{-7}&7.4584\times 10^{-1}&2.6980\times 10^{-5}\\ 17&2.6826\times 10^{-1}&8.3207\times 10^{-1}&5.5187\times 10^{-7}&7.2919\times 10^{-1}&5.8263\times 10^{-5}\\ \infty&2.1746\times 10^{-1}&8.2380\times 10^{-1}&1.2025\times 10^{-6}&7.0683\times 10^{-1}&1.1793\times 10^{-4}\\ \hline\cr\eta&m_{46}&m_{47}&m_{48}\\ 5&0&0&\frac{24}{5}\\ 7&1.2222\times 10^{-1}&2.4505\times 10^{-1}&4.7008\\ 10&2.0315\times 10^{-1}&4.0914\times 10^{-1}&4.6366\\ 17&2.7309\times 10^{-1}&5.5322\times 10^{-1}&4.5779\\ \infty&3.6056\times 10^{-1}&7.3790\times 10^{-1}&4.4916\\ \hline\cr\hline\cr\end{array}$